OpenStudy (anonymous):

All identities are equations, and all equations are identities. True or false?

OpenStudy (anonymous):

\[x+2=7\]

OpenStudy (anonymous):

I think true, since both share the "equality" condition?

OpenStudy (anonymous):

An identity is true for all values of x. An equation is true for only certain values of x.

OpenStudy (anonymous):

\[x^2=x\times x\] This is an identity Equations are true for only a set of values \[x^2+2x+1=0\] True for only x=-1

OpenStudy (anonymous):

So it is false?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

not all equations are identities

OpenStudy (anonymous):

'Kay, gotcha. Thanks, guys!